Metlock Company is negotiating to lease a piece of equipment to MTBA, Inc. MTBA requests that the lease be for 9 years. The equipment has a useful life of 10 years. Metlock wants a guarantee that the residual value of the equipment at the end of the lease is at least $4,000. MTBA agrees to guarantee a residual value of this amount though it expects the residual value of the equipment to be only $2,000 at the end of the lease term. If the fair value of the equipment at lease commencement is $105,000, what would be the amount of the annual rental payments Metlock demands of MTBA, assuming each payment will be made at the beginning of each year and Metlock wishes to earn a rate of return on the lease of 10%?

Got 17938 (rounded) wrong
Answer #1

Present value of residual value = 4000*0.42410 = 1696.40

PV factor of $1 @ i = 10%, n =9 is 0.42410

Amount to be recovered through lease payments = fair value - Present value of residual value = 105000-1696.40 = $103303.60

Annual rental payments = 103303.60/6.33493 = $16307

PV annuity due factor of $1 @ i = 10%, n =9 is 6.33493